# The heat flow on glued spaces with varying dimension

### v1

@@ 3.3 Compact embedding of $H^{1}_{*}$
We prove the compactness of the embedding through sequentially compactness and an inner approximation up to the boundary. The idea of this proof is based on the techniques of [55, Chapter 5]. For the convenience of the reader, we explain the details in our case here.
@@
Let $(X,d,\mu)$ be as in Theorem 2.2 and $\operatorname*{\mathcal{E}}$ as before. Then, $H^{1}_{*}$ is compactly embedded into $L^{2}(X,\mu)$, i.e. for $(u_{n})_{n\in\mathbb{N}}\subset H^{1}_{*}$ bounded, there exists a subsequence that converges strongly in $L^{2}(X,\mu)$.
@@
and define $X_{\varepsilon}:=\{x\in X:\operatorname*{dist}(x,\partial X)<\varepsilon\}$ where $\partial X:=\bigcup\partial M_{i}$. We will prove that this sequence has a converging subsequence in $L^{2}$. Choose a open covering of $X\setminus X_{\varepsilon}$ by balls, via doing this for both $M_{i}$ and taking the union. This covering consists of $q$ balls of fixed radius $r<\varepsilon$ s.t.
@@
By Banach-Alaoglu there exists a subsequence and $u\in L^{2}$ s.t. $u_{n}\rightharpoonup u$ in $L^{2}$. Let $\omega_{n,m}:=u_{n}-u_{m}$ be the difference. We aim to show $\omega_{n,m}\to 0$ in $L^{2}(X\setminus X_{\varepsilon})$.
With $(\omega_{n,m})_{i}$ as the mean value, we can estimate
@@
[…]
The first part I can be estimated by the Poincare inequality on balls on $M_{i}$, see Section 2:
@@
The second part II directly yields:
@@
Since $\omega_{n,m}\rightharpoonup 0$ in $L^{2}$ as $n,m\to\infty$ we have $\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\to 0$. Now, fix $\delta>0$ and plug in $r=\sqrt{\frac{\delta}{l}}$:
@@
This holds for $n,m$ large enough since $q$ is bounded. Thus, we can find $n_{*},m_{*}$ dependent on $\delta$ s.t. for $n>n_{*},m>m_{*}$ this inequality holds. We conclude that $\omega_{n,m}$ converges to 00 in $L^{2}$.
@@
$\Gamma_{X}(\varepsilon)$ $:=\sup\limits_{\left\lVert u\right\rVert_{H^{1}_{*}(X)}=1}\int\limits_{X_{\varepsilon}}|u|^{2}\mathop{}\!\mathrm{d}\mu,$
$\Gamma_{X}(0)$ $:=\lim\limits_{\varepsilon\searrow 0}\Gamma_{X}(\varepsilon)$
Since the manifolds are smooth with smooth boundary, we have $\Gamma_{X}(0)=0$.
Take $u_{n}$ a sequence in $L^{2}(X)$ s.t. $\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}\leq 1\;\forall n$. Then, for every $\delta>0$ we find $\varepsilon$ s.t. $\Gamma_{X}(\varepsilon)<\delta$ and a subsequence $u_{n}$ s.t.
@@
$\forall n\geq n_{0}(\varepsilon)$ $:\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}\leq\delta.$
The sequence is constructed via a diagonal sequence. First, take a subsequence $u_{n}\rightharpoonup u$ in $L^{2}(X)$. Afterwards, for a sequence $\varepsilon\to 0$ we can find refining subsequences (the existence is given by a diagonal sequence) s.t. $u_{n}\to u$ in $L^{2}(X\setminus X_{\varepsilon})$ and for $n\geq n_{0}(\varepsilon)$:
$\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}\leq\delta.$
@@
$\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X)}^{2}$ $=\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X\setminus X_{\varepsilon})}^{2}+\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon})}^{2}+\left\lVert u\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon})}^{2}$
$\lesssim\delta+\delta\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}^{2}\lesssim\delta$
We used the lower semi-continuity of the $L^{2}$ norm w.r.t. to the weak $L^{2}$ convergence. With $\delta\to 0$ this implies the compact embedding. ∎
@@
Now, we can combine these calculations and conclude that only constants lie in the kernel of $A$.
Given the assumptions above, $H^{1}_{*}$ has a continuity condition. Thus, assume $f\in\text{Ker}(A)$. Then
$0=-(Af,f)=\operatorname*{\mathcal{E}}(f).$
On each manifold, we fix charts and take partition of unities on the charts which by the transformation to $\mathbb{R}^{n}$ yield
@@
Since $\omega_{i}>0$, we conclude that $f$ is constant a.e. on each chart. By approximation and since the manifolds are connected, $f$ is constant on each manifold. Now, we can infer the value of $f$ on $L$ q.e. by the mean-value on each manifold and it has to be the same on both sides (since the capacities are equivalent), which yields that $f$ has to be constant on $X$, i.e. the kernel of $A$ are the constant functions. Together with the facts that $A$ is a selfadjoint, non-positive operator and has a compact resolvent by operator-theory, it follows that the spectrum of $A$ is discrete with no finite accumulation point and $A$ has a spectral gap, c.f. [64, Chapter 8]. This is equivalent to satisfying a global Poincare type inequality on the orthogonal complement of the constants $\langle 1\rangle^{\perp}$. To see this, we can write down the condition on the orthogonal complement of the kernel. There exists $\lambda\geq 0$ s.t.
$\lambda\left\lVert f-f_{X}\right\rVert_{L^{2}(X,\mu)}\leq-(Af,f)=\mathcal{E}(f).$
Next, we apply the spectral theorem which represents $A$ as a multiplication operator $a\leq 0$ on $L^{2}$ with $a\leq-\lambda$ on the orthogonal complement of the constants. For f with $\int f\mathop{}\!\mathrm{d}\mu=0$ this yields
$\left\lVert S_{h}f\right\rVert_{L^{2}(\mu)}\leq e^{-\lambda h}\left\lVert f_{0}\right\rVert_{L^{2}(\mu)}\to 0.$
Therefore, solutions converge exponentially fast to the constants that are their mean value.
This, in turn, proves the irreducibility or ergodicity of $S_{h}$ following the Definitions 2.7 and 2.8.
@@
By monotonicity we know that if $A\subseteq B$ then
@@
This holds as $T_{t}$ is monotone since $S_{h}$ is linear and monotone (positive). By [20, Theorem 1.4.1], $S_{h}$ is Markovian as it is the semigroup associated to a Dirichlet form, i.e. if $0\leq f\leq 1$ then $0\leq S_{h}\leq 1$. Thus, by linearity of $S_{h}$ it is positivity preserving on a.e. bounded functions and therefore monotone.
By the upper bound ($S_{t}\mathds{1}_{A}\leq 1$ a.e.), $S_{s}\mathds{1}_{A}\leq T_{s}(A)$ and thus
@@
If $\mu(E)>0$ we have $T_{t}(E)\to X$ as $t\nearrow\infty$ by the observation that all functions converge to their mean as a constant function under the flow. If $T_{t}(E)\subseteq E$ for $t>0$ then
@@
To prove irreducibility, we can plug in $f\equiv 1$ and get $S_{h}(\mathds{1}_{E})=\mathds{1}_{E}$ $\mu-$a.e. which implies $T_{h}(E)=E$ a.e.. ∎
@@ 4 Generalizations
In this section, we show how the proof and techniques of Section 3 can be generalized to the case of more then two connected manifolds. Moreover, we give an idea how to incorporate more general weights.
We still rely on the equivalence of the capacities to get a trace in the capacity of our Sobolev space. Thus, the effective dimension (taking into account the weight) of the manifolds at their intersection cannot vary too much. Heuristically spoken, the difference of $k$ and the effective dimension minus one has to be less then one.
For general weights $\omega_{1},\omega_{2}\in A_{2}$, we can follow the same calculations to get lower bounds on the capacities:
@@
$\quad\left.+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
@@
$\quad\left.+2(n_{2}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Additionally, if $\omega_{2}$ satisfies the $L-$Muckenhoupt type condition on the intersection,
$\int_{L_{r}}\omega_{2}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\int_{L_{r}}\frac{1}{\omega_{2}}\mathop{}\!\mathrm{d}\text{vol}_{M_{2}}\simeq r^{2(n_{2}-k)},$
we can simplify these expressions to
$\text{Cap}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
$\text{Cap}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Example.
We consider a submanifold $L$ in $M$ with $n:=\dim M$ and $k:=\dim L\neq n-1$ with weight $\omega$ which locally around $L$ has the form
$\omega(x):=\frac{1}{\operatorname*{dist}(x,L)^{\alpha}}.$
Then locally,
$\frac{\mathcal{L}^{n}}{\omega}=\int\limits_{L_{r}}\frac{1}{\omega}\mathop{}\!\mathrm{d}x$$\simeq\begin{cases}\frac{r^{n-k+\alpha}}{n-k+\alpha}&\text{if }\alpha>-(n-k),\\ \infty&\text{else.}\end{cases}$
$\int\limits_{0}^{R}\rho\mu_{i}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho$ $=\int\limits_{0}^{R}\rho^{1-n+k+\alpha}\mathop{}\!\mathrm{d}\rho<\infty,$
i.e. $\alpha>n-k-2$ which coincides with our intuition that the difference of $k$ and effective dimension minus one, $n-\alpha-1$, has to be less than one.
Moreover, we can prove an upper capacity bound which shows that our lower bound is optimal up to a constant. This upper bound of the capacity can be obtained through an explicit competitor. First, consider $L_{r}$ in $L_{2r}$ and the linear competitor $u(x)=\left(1-\frac{\operatorname*{dist}(x,L_{r})}{r}\right)\mathds{1}_{L_{2r}}\in\operatorname*{Lip}\cap C_{c}(L_{2r})$:
@@
$\leq\int\limits_{L_{2r}}r^{-2}\mathop{}\!\mathrm{d}\mu_{i}$
@@
This can be used iteratively to compute the capacity. With $k\in\mathbb{N}$ s.t. $2^{k-1}r\leq R\leq 2^{k}r$ we get:
$\text{Cap}_{2,\mu_{i}}(L_{r},L_{R})$ $\lesssim\text{Cap}_{2,\mu_{i}}(L_{r},L_{2^{k}}r)$
$\leq\left(\sum\limits_{j=0}^{k-1}\text{Cap}_{2,\mu_{i}}(L_{2^{j}r},L_{2^{j+1}r})^{-1}\right)^{-1}$
@@
$\leq\left(\int\limits_{L_{2^{k}r}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}}{\omega_{i}}\right)^{-1}$
@@
$\leq\left(\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{i}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{i}}}{\omega_{i}}\right)^{-1}.$
@@
The proof of the first inequality is analogous to [36, Lemma 2.16]. With the simplification from Section 3,
@@
$\text{Cap}_{2,\mu_{i}}(L,L_{R})\lesssim\left(R^{-2(n_{i}-k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(L_{R})+2(n_{i}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{i}-2k-1)}\frac{\mathcal{L}^{n_{i}}}{\omega_{i}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$

### v2

@@ 3.3 Compact Embedding of $H^{1}_{*}$
We prove the compactness of the embedding through sequentially compactness and an inner approximation up to the boundary. The idea of this proof is based on the techniques of [61, Chapter 5]. For the convenience of the reader, we explain the details in our case here.
@@
[…]
@@
and define $X_{\varepsilon}\coloneqq\{x\in X:\operatorname*{dist}(x,\partial X)<\varepsilon\}$ where $\partial X\coloneqq\bigcup\partial M_{i}$. We prove that this sequence has a converging subsequence in $L^{2}$. We choose an open covering of $X\setminus X_{\varepsilon}$ by balls by doing this for both $M_{i}$ individually and taking the union. For this we follow [61, Lemma 5.2]. This covering consists of $q$ balls of fixed radius $r<\varepsilon$ such that
@@
By Banach-Alaoglu there exists a subsequence and $u\in L^{2}$ such that $u_{n}\rightharpoonup u$ in $L^{2}$. Let $\omega_{n,m}\coloneqq u_{n}-u_{m}$ be the difference. We aim to show $\omega_{n,m}\to 0$ in $L^{2}(X\setminus X_{\varepsilon})$.
With $(\omega_{n,m})_{i}$ as the mean value in $B_{i}$, we can estimate
@@
$=:\text{I}+\text{II}.$
The first part I can be estimated by the Poincaré inequality on balls on $M_{i}$, see Definition 2.9 and Subsection 2.1:
@@
The second part II directly yields (with a possibly larger $\nu$ that only depends on the doubling constant of $\text{vol}_{M_{i}}$):
@@
Since $\omega_{n,m}\rightharpoonup 0$ in $L^{2}$ as $n,m\to\infty$, we have $\int\limits_{B_{i}}\omega_{n,m}\mathop{}\!\mathrm{d}\mu\to 0$. Now, fix $\delta>0$ and plug in $r=\sqrt{\frac{\delta}{l}}$:
@@
This holds for $n,m$ large enough since $q$ is bounded independently of $n,m$. Thus, we can find $n_{*},m_{*}$ dependent on $\delta$ such that for $n>n_{*},m>m_{*}$ this inequality holds. We conclude that $\omega_{n,m}$ and thus $u_{n}$ converges to 0 strongly in $L^{2}$.
@@
$\Gamma_{X}(\varepsilon)$ $\coloneqq\sup\limits_{\left\lVert u\right\rVert_{H^{1}_{*}(X)}=1}\int\limits_{X_{\varepsilon}}|u|^{2}\mathop{}\!\mathrm{d}\mu,$
$\Gamma_{X}(0)$ $\coloneqq\lim\limits_{\varepsilon\searrow 0}\Gamma_{X}(\varepsilon).$
Since the manifolds are smooth with smooth boundary and bounded weights in a neighborhood of the boundary, we have $\Gamma_{X}(0)=0$, see [61, Chapter 5].
Take $u_{n}$ a sequence in $L^{2}(X)$ such that $\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}\leq 1\;\forall n$. Then, for a sequence $\delta_{k}>0$, we find $\varepsilon_{k}$ such that $\Gamma_{X}(\varepsilon_{k})<\delta_{k}$ and a subsequence $u_{n}$ such that
@@
$\forall n\geq n_{0}(k)$ $:\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon_{k}})}^{2}\leq\delta_{k}.$
The sequence is constructed via a diagonal sequence. First, take a subsequence $u_{n}\rightharpoonup u$ in $L^{2}(X)$. Afterwards, for $\varepsilon_{k}$, we can find refining subsequences such that $u_{n}\to u$ in $L^{2}(X\setminus X_{\varepsilon_{k}})$ and for $n\geq n_{0}(k)$:
$\left\lVert u-u_{n}\right\rVert_{L^{2}(X\setminus X_{\varepsilon_{k}})}^{2}\leq\delta_{k}.$
The existence of the whole sequence is given as a diagonal sequence.
@@
$\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X)}^{2}$ $=\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X\setminus X_{\varepsilon_{k}})}^{2}+\limsup_{n\to 0}\left\lVert u_{n}-u\right\rVert_{L^{2}(X_{\varepsilon_{k}})}^{2}$
$\lesssim\delta_{k}+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon_{k}})}^{2}+\left\lVert u\right\rVert_{L^{2}(X_{\varepsilon_{k}})}^{2}$
$\lesssim\delta_{k}+\limsup_{n\to 0}\left\lVert u_{n}\right\rVert_{L^{2}(X_{\varepsilon_{k}})}^{2}$
$\lesssim\delta_{k}+\delta_{k}\left\lVert u_{n}\right\rVert_{H^{1}_{*}(X)}^{2}\lesssim\delta_{k}.$
We used the lower semi-continuity of the $L^{2}$-norm with respect to the weak $L^{2}$-convergence. With $\delta_{k}\to 0$, this implies the compact embedding. ∎
@@
Now, we can combine these calculations and conclude that only constants lie in the kernel of $\operatorname*{\mathcal{A}}$.
Given the above assumptions, $H^{1}_{*}$ has a continuity or trace condition, see Remark 3. Thus, assume $f\in\text{Ker}(\operatorname*{\mathcal{A}})$. Then,
$0=-(\operatorname*{\mathcal{A}}f,f)=\operatorname*{\mathcal{E}}(f).$
On each manifold, we fix charts and take partition of unities on the charts which by the transformation to $\mathbb{R}^{n_{i}}$ yield
@@
Since $\omega_{i}>0$, we conclude that $f$ is constant a.e. on each chart. By approximation and since the manifolds are connected, $f$ is constant on each manifold. Now, we can infer the value of $f$ on $L$ q.e. by the mean value on each manifold and it has to be the same on both sides (since the capacities are equivalent), which shows that $f$ has to be constant in $X$, that is, the kernel of $\operatorname*{\mathcal{A}}$ consists precisely of constant functions.
Together with the facts that $\operatorname*{\mathcal{A}}$ is a selfadjoint, nonpositive operator and has a compact resolvent by operator-theory, it follows that the spectrum of $\operatorname*{\mathcal{A}}$ is discrete with no finite accumulation point and $\operatorname*{\mathcal{A}}$ has a spectral gap and an orthonormal basis of eigenfunctions, cf. [69, Proposition 8.8]. This is equivalent to satisfying a global Poincaré type inequality on the orthogonal complement of the constants $\langle 1\rangle^{\perp}$. To see this, we can write down the condition on the orthogonal complement of the kernel. There exists $\lambda>0$ such that
$\lambda\left\lVert f-f_{X}\right\rVert_{L^{2}(X,\mu)}\leq-(\operatorname*{\mathcal{A}}f,f)=\mathcal{E}(f).$
Next, we apply the spectral theorem which represents $\operatorname*{\mathcal{A}}$ as a multiplication operator $a\leq 0$ on $L^{2}(X)$ with $a\leq-\lambda$ on the orthogonal complement of the constants. For $f$ with $\int f\mathop{}\!\mathrm{d}\mu=0$, this yields
$\left\lVert S_{t}f\right\rVert_{L^{2}(\mu)}\leq e^{-\lambda t}\left\lVert f_{0}\right\rVert_{L^{2}(\mu)}\to 0$
as $t\to\infty$. Therefore, solutions converge exponentially fast to the constants that are their mean values.
This, in turn, proves the irreducibility or ergodicity of $S_{t}$ following the Definitions 2.11 and 2.12:
@@
By monotonicity, we know that if $A\subseteq B$ then
@@
This holds as $T_{t}$ is monotone since $S_{t}$ is linear and monotone (positive) as shown now. By [24, Theorem 1.4.1], $S_{t}$ is Markovian, as it is the semigroup associated with a Dirichlet form, that is, if $0\leq f\leq 1$ then $0\leq S_{h}f\leq 1$. Thus, by linearity of $S_{t}$, it is positivity preserving on bounded function and therefore monotone.
By the upper bound ($S_{t}\chi_{A}\leq 1$ a.e.), $S_{s}\chi_{A}\leq T_{s}(A)$ and thus
@@
If $\mu(E)>0$, we have $T_{t}(E)\to X$ as $t\nearrow\infty$ by the observation that all functions converge to their mean as a constant function under the flow. If $T_{t}(E)\subseteq E$ for $t>0$ then
@@
To prove irreducibility, we can plug in $f\equiv 1$ and get $S_{h}(\chi_{E})=\chi_{E}$ $\mu$-a.e. which implies $T_{h}(E)=E$ a.e. ∎
@@ 4 Generalizations
In this section, we show how the proof and techniques of Section 3 can be generalized to the case of more than two connected manifolds. Moreover, we give an idea on how to incorporate more general weights.
Theorem 4.1 (Heat flow on the glued manifold II).
Let $(X,d,\mu)$ be a metric space with $X=\bigcup\limits_{i=1}^{m}M_{i}$ being a union of smooth, bounded, compact, connected, Riemannian manifolds whose intersection $L_{ij}\coloneqq M_{i}\cap M_{j}$ are disjoint, transversal, nonempty and lie in the inner of both manifolds. Equip $X$ with the glued distance $d$ and assume that $X$ is path-wise connected. Furthermore, let
$\mu\coloneqq\sum\limits_{i=1}^{m}\omega_{i}\cdot\text{vol}_{M_{i}}$
fulfill $\mu(L_{ij})=0$ and be locally $N_{ij}$-doubling.
Moreover, assume nonnegative weights $\omega_{i}\in A_{2}(M_{i}),\omega_{i}>0$ such that they are bounded in a neighborhood of $\partial M_{i}$ and on each intersection $L\in\{L_{ij}\}$ the following holds for one of both manifolds:
  • $\mu_{i}$ satisfies the $L_{ij}$-Muckenhoupt type condition, see Definition 3.2,
  • $\int\limits_{0}^{R}\frac{N_{ij}(\rho)\rho}{\mu_{i}(L_{\rho})}\mathop{}\!\mathrm{d}\rho,\int\limits_{0}^{R}\frac{\rho}{\mu_{i}(L_{\rho})}\mathop{}\!\mathrm{d}\rho<\infty$.
Then, the canonical heat flow is ergodic and irreducible.
Remark.
In Example 4.2, we give a class of measures that satisfy the $L$-Muckenhoupt condition. The other conditions are required for the capacity to be positive, and thus for the connectedness of the manifolds. These can be read intuitively as the fact that the effective codimension of the intersection must be strictly between 0 and $2$ and that the jump in the effective dimension cannot be too large.
As in the case of two manifolds, we still rely on the equivalence of the capacities to get a trace in our Sobolev space. Thus, the effective dimension (taking into account the weight) of the manifolds at their intersection cannot vary too much. In terms of heuristically speaking, the difference of $k$ and the effective dimension minus one must be less than one. We first prove the case of two manifolds with general weights, as in Theorem 4.1, intersecting, and then discuss that all the proofs carry over to the case of disjoint intersections.
We begin with lower bounds. For general weights $\omega_{1},\omega_{2}\in A_{2}$ on $M_{i}$ with $\text{dim}(M_{i})=n_{i}$, we can follow the same calculations as above to obtain lower bounds on the capacities:
@@
$\qquad\left.+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1},$
@@
$\qquad\left.+2(n_{2}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Additionally, if $\omega_{2}$ satisfies the $L$-Muckenhoupt type condition on the intersection, see Definition 3.2, we can simplify these expressions:
${\text{Cap}}_{2,\mu_{1}}(L,L_{R})$ $\gtrsim\left(N(R)R^{2}\mu_{2}(L_{R})^{-1}+2(n_{1}-k-1)\int\limits_{0}^{R}N(\rho)\rho\mu_{2}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho\right)^{-1},$
${\text{Cap}}_{2,\mu_{2}}(L,L_{R})$ $\gtrsim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Moreover, we can prove an upper capacity bound, which shows that our lower bounds are optimal up to a constant. This upper bound on the capacity can be obtained through an explicit competitor. First, consider $L_{r}$ in $L_{2r}$ and the linear competitor $u(x)=\left(1-\frac{\operatorname*{dist}(x,L_{r})}{r}\right)\chi_{L_{2r}}\in\operatorname*{Lip}\cap C_{c}(L_{2r})$:
@@
$\leq\int\limits_{L_{2r}}r^{-2}\mathop{}\!\mathrm{d}\mu_{2}$
@@
This can be used iteratively to compute the capacity. With $l\in\mathbb{N}$ such that $2^{l-1}r\leq R\leq 2^{l}r$, we get:
${\text{Cap}}_{2,\mu_{2}}(L_{r},L_{R})$ $\lesssim{\text{Cap}}_{2,\mu_{2}}(L_{r},L_{2^{l}r})$
$\leq\left(\sum\limits_{j=0}^{l-1}{\text{Cap}}_{2,\mu_{2}}(L_{2^{j}r},L_{2^{j+1}r})^{-1}\right)^{-1}$
@@
$\leq\left(\int\limits_{L_{2^{l}r}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{2}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{2}}}{\omega_{2}}\right)^{-1}$
@@
$\leq\left(\int\limits_{L_{R}\setminus L_{r}}\frac{1}{\operatorname*{dist}(y,L)^{2(n_{2}-k-1)}}\frac{\mathop{}\!\mathrm{d}\mathcal{L}^{n_{2}}}{\omega_{2}}\right)^{-1}.$
@@
The proof of the first inequality is analogous to [40, Lemma 2.16]. With the simplification of Section 3,
${\text{Cap}}_{2,\mu_{2}}(L,L_{R})$ $\lesssim\left(R^{-2(n_{2}-k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{R})\right$.
@@
$\qquad\qquad\left.+2(n_{2}-k-1)\int\limits_{0}^{R}\rho^{-(2n_{2}-2k-1)}\frac{\mathcal{L}^{n_{2}}}{\omega_{2}}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$